OpenStudy (anonymous):

Simplify the expression involving rational exponents: (-27/125)^(-2/3)

OpenStudy (anonymous):

Start with cuberoot( (-27/125)^(-2) )... and then ? pleaseandthanks

OpenStudy (mertsj):

\[(\sqrt[3]{\frac{-27}{125}})^2=(\frac{\sqrt[3]{-27}}{\sqrt[3]{125}})^2=(\frac{-3}{5})^2=\frac{9}{25}\]

OpenStudy (anonymous):

Wrong. 25/9

OpenStudy (mertsj):

Yes. I missed that the exponent is negative so your first step should be to flip the fraction. Sorry.